OpenStudy (anonymous):

4. Can a spherical thank with a 50ft diameter hold enough water to fill a swimming pool that has a 50m length, a 25m width, and a constant 2m depth? Explain your answer and provide work and calculations to support your response.

OpenStudy (anonymous):

volume of a sphere = (4/3)*pi*r^3 = (4/3)*pi*(50/2)^3 = 65449.8 ft^3 volume of pool = 50*25*2 = 2500 ft^3 so yes. quite a bit more.